# In-and-Out: Algorithmic Diffusion for Sampling Convex Bodies

### v1

@@ 1.3 Discussion
Here we make a few remarks, contrasting our results with known ones.
@@
Previous uniform samplers like the $\mathsf{Ball\ walk}$ are made lazy (i.e., with probability $1/2$, the Markov chain does nothing), to ensure convergence to the target stationary distribution. However, our algorithm does not need this, since we directly show that our sampler contracts towards a uniform distribution.
@@
We remark that these two different bounds also place the previously known mixing guarantees for $\mathsf{Ball\ walk},\mathsf{Speedy\ walk}$ in a unified framework. Existing tight guarantees for $\mathsf{Speedy\ walk}$ are in TV distance and based on the log-Sobolev constant, assuming an oracle for implementing each step [lee2017eldan]. The known convergence guarantees of $\mathsf{Ball\ walk}$ (see Appendix B for details), namely the mixing time of $\widetilde{\mathcal{O}}(Md^{2}D^{2}\log\frac{1}{\varepsilon})$ for TV distance, are for the composite algorithm [$\mathsf{Speedy\ walk}+$rejection sampling]. Here $\mathsf{Speedy\ walk}$ records only the accepted steps of $\mathsf{Ball\ walk}$, so its stationary distribution differs slightly from the uniform distribution (and can be corrected with a post-processing step). On the other hand, $\mathsf{In\text{-}and\text{-}Out}$ actually converges to $\pi^{\mathcal{K}}$ without any adjustments and achieves stronger Rényi divergence bounds in the same asymptotic complexity. Our analysis shows that the mixing guarantee is determined by isoperimetric constants of the target (Poincaré or log-Sobolev).
@@
The $\mathsf{Ball\ walk}$’s largest possible step size is of order $1/\sqrt{d}$ (see Appendix B) to keep the rejection probability bounded by a constant. This bound could also be viewed as an “effective” step size of $\mathsf{In\text{-}and\text{-}Out}$. This follows from the fact that the $\ell_{2}$-norm of the Gaussian $\mathcal{N}(0,hI)$ is concentrated around $\sqrt{hd}$, and we will set the variance $h$ of $\mathsf{In\text{-}and\text{-}Out}$ to $\widetilde{\mathcal{O}}(1/d^{2})$, so we have $\sqrt{hd}\approx 1/\sqrt{d}$.
@@
How can we compare $\mathsf{In\text{-}and\text{-}Out}$ with $\mathsf{Speedy\ walk}?$ Iterating speedy steps leads to a biased distribution (one that is proportional to the local conductance). As clarified in Remark 2, one step of (a Gaussian version of) $\mathsf{Speedy\ walk}$ can be understood as a step of backward heat flow. Therefore, if one can control the isoperimetric constants of the biased distribution along the trajectory of the backward flow, then contraction of $\mathsf{Speedy\ walk}$ toward the biased distribution will follow from the simultaneous backward analysis.
@@ 2 Preliminaries
$\|\mu-\nu\|_{\mathsf{TV}}:=\sup_{B\in\mathcal{F}}\lvert\mu(B)-\nu(B)\rvert\,,$
@@
$\mathcal{W}_{2}^{2}(\mu,\nu):=\inf_{\gamma\in\Gamma(\mu,\nu)}\mathbb{E}_{(X,Y)\sim\gamma}[\|X-Y\|^{2}]\,,$
@@
We recall two important functional inequalities of a distribution. We use $\nu$ to denote a probability measure over $\mathbb{R}^{d}$.
@@
We say that $\nu$ satisfies a Poincaré inequality (PI) with parameter $C_{\mathsf{PI}}(\nu)$ if for all smooth functions $f:\mathbb{R}^{d}\to\mathbb{R}$,
$\mathrm{Var}_{\nu}f\leq C_{\mathsf{PI}}(\nu)\,\mathbb{E}_{\nu}[\|\nabla f\|^{2}]\,,$
where $\mathrm{Var}_{\nu}f\coloneqq\mathbb{E}_{\nu}\lvert f-\mathbb{E}_{\nu}f\rvert^{2}$.
@@
We say that $\nu$ satisfies a log-Sobolev inequality (LSI) with parameter $C_{\mathsf{LSI}}(\nu)$ if for all smooth functions $f:\mathbb{R}^{d}\to\mathbb{R}$,
$\mathsf{Ent}_{\nu}(f^{2})\leq 2C_{\mathsf{LSI}}(\nu)\,\mathbb{E}_{\nu}[\|\nabla f\|^{2}]\,,$
@@
where $\mathsf{FI}(\mu\mathbin{\|}\nu):=\mathbb{E}_{\mu}[\|\nabla\log\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\|^{2}]$ is the Fisher information of $\mu$ with respect to $\nu$.
We recall the data-processing inequality for Rényi divergence and $f$-divergence.
@@
The aforementioned functional inequalities allow us to show exponential contraction of various divergences, through the following helpful inequality.
@@ 3 Analysis
One successful iteration of $\mathsf{In\text{-}and\text{-}Out}$ is exactly the same as the composition of running the forward heat flow and then backward heat flow, both for time $h$. Its validity is justified in Section 1.1, and we give its proof below.
@@
Regarding the backward part, it is known from [chen2022improved, Lemma 14] that the construction of the time-reversal SDE ensures that $(Z_{h}^{\leftarrow},Z_{0}^{\leftarrow})$ and $(Z_{0},Z_{h})$ have the same joint distribution, when $Z_{0}\sim\pi^{X}$ (and so $Z_{h}\sim\pi^{Y}$). Hence, $\mathsf{law}(Z_{h}^{\leftarrow}|Z_{0}^{\leftarrow}=y)=\mathsf{law}(Z_{0}|Z_{h}=y)=\pi^{X|Y=y}$, where the last equality follows from $(Z_{0},Z_{h})\sim\pi$. Since we initialize ($\mathsf{BH}$) with $Z_{0}^{\leftarrow}=y\sim\mu_{k+1}^{Y}$, we have
@@
For (1), we collect in Section 3.1 some important implications of functional inequalities, e.g. the Poincaré and log-Sobolev inequalities, for the uniform distribution. Then in Section 3.2, we exploit the flow perspective of the algorithm to obtain the mixing guarantees. To this end, we revisit the proofs for the contraction results of forward and backward heat flows, and then extend them to our constrained setting:
Theorem 13 (Adapted for uniform distributions, [chen2022improved]).
@@
This reduces the problem of obtaining a mixing guarantee to that of demonstrating a functional inequality on the target distribution. For this, it is not strictly necessary that $\mathcal{K}$ be convex.
On the other hand, convexity of $\mathcal{K}$ is crucial for the proof of (2). We show in Section 3.3 that the failure probability remains under control by taking a suitable variance $h$ and threshold $N$. We then show that the expected number of trials per iteration is of order $\log N$, not $N$:
@@
Let $\mathcal{K}$ be any convex body in $\mathbb{R}^{d}$ presented by a well-defined membership oracle, $\pi^{X}$ the uniform distribution over $\mathcal{K}$, and $\mu$ an $M$-warm initial distribution with respect to $\pi^{X}$. For any given $m\in\mathbb{N}$ and $\eta\in(0,1)$, set $Z=\frac{9mM}{\eta}(\geq 9)$, $h=\frac{\log\log Z}{2d^{2}\log Z}$ and $N=Z(\log Z)^{4}=\widetilde{\mathcal{O}}(\frac{mM}{\eta})$. Then, the failure probability of one iteration of $\mathsf{In\text{-}and\text{-}Out}$ is at most $\eta/m$. Moreover, the expected number of membership queries needed per iteration is $\mathcal{O}\bigl{(}M(\log\frac{mM}{\eta})^{4}\bigr{)}$.
@@ 3.1 Functional inequalities
The contraction of an outer loop of our algorithm is controlled by isoperimetry of the uniform distribution $\pi^{X}$, which is described precisely by a functional inequality. The most natural ones to consider in this setting are the Poincaré inequality (7) and log-Sobolev inequality (8). In Appendix C, we provide a more detailed discussion of how these are related to other important notions of isoperimetry, such as the Cheeger and log-Cheeger inequalities.
@@
Lemma 16 ([vempala2019rapid, Lemma 9]).
@@
where $\mathsf{RF}_{q}(\mu\mathbin{\|}\nu):=q\,\mathbb{E}_{\nu}\bigl{[}\bigl{(}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\bigr{)}^{q}\|\nabla\log\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\|^{2}\bigr{]}\,/\,\mathbb{E}_{\nu}\bigl{[}\bigl{(}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\bigr{)}^{q}\bigr{]}$ is the Rényi Fisher information of order $q$ of $\mu$ with respect to $\nu$.
The log-Sobolev inequality paired with the KL-divergence (8) can be understood as a special case of the following inequality [^2] Such inequalities are often called Polyak-Łojasiewicz inequalities, which say for $f:\mathbb{R}^{d}\to\mathbb{R}$, and all $y\in\mathbb{R}^{d}$ that $f(y)\leq c\,\|\nabla f(y)\|^{2}$ for some constant $c$, if $\min f(x)=0$. paired with the $q$-Rényi divergence for $q\geq 1$.
Lemma 17 ([vempala2019rapid, Lemma 5]).
@@
We have collected below the functional inequalities used to establish the mixing of our algorithm (see Appendix C for a detailed presentation).
@@
Let $\mathcal{K}\subset\mathbb{R}^{d}$ be a convex body with diameter $D$, and $\pi$ be the uniform distribution over $\mathcal{K}$. Then, $C_{\mathsf{PI}}(\pi)\lesssim\|\mathrm{Cov}(\pi)\|_{\mathsf{op}}\,\log d$ and $C_{\mathsf{LSI}}(\pi)\lesssim D^{2}$. If $\pi$ is isotropic, then $C_{\mathsf{PI}}(\pi)\lesssim\log d$ and $C_{\mathsf{LSI}}(\pi)\lesssim D$.
@@ 3.2 Contraction and mixing
We start by analyzing how many outer iterations of $\mathsf{In\text{-}and\text{-}Out}$ are required to be $\varepsilon$-close to $\pi^{X}$, the uniform distribution over $\mathcal{K}$. The contraction of Algorithm 1 comes from analyzing Lines 2 and 3 through the perspective of heat flows (see Section 1.1). To exploit this view, we first revisit the previous contraction analysis in [chen2022improved], which is carried out for distributions with smooth densities. Although the uniform distribution is not even continuous, we prove a technical lemma (Lemma 22) that enables us to extend previously known results to the uniform distribution. Lastly, combining the previous results with our technical lemma, we obtain clean contraction results of Algorithm 1 toward the uniform distribution $\pi^{X}$ in Theorem 23.
@@
In this part, we review the contraction results for heat flow and its time-reversal [chen2022improved], which are intimately connected with our algorithm. We also provide key technical ingredients needed for its proof, such as the computations for measures evolving under simultaneous forward/backward heat flows. We refer interested readers to Appendix D for additional details. Only in Part I, we assume that $\nu$ denotes a probability measure with smooth density.
@@
Although there are many ways to define a “reversal” of $P_{h}$, we will use the notion of adjoint introduced by [klartag2021spectral], which is the most immediately useful.
@@
Instead of considering the change in metrics along the evolution of $\mu P_{t}$ with respect to “fixed” $\nu$, it will be useful to consider the simultaneous evolution of $\mu P_{t},\nu P_{t}$ (and similarly $\mu Q_{t},(\nu P_{h})Q_{t}$). This type of computation was carried out for specific metrics in earlier work [vempala2019rapid, chen2022improved]. The following is a more generalized form of one appearing in [yuan2023class, Lemma 2]. In the lemma below, we consider an arbitrary diffusion equation with corresponding Fokker-Planck equation:
@@
$\int\bigl{[}-\nabla\cdot(b_{t}\nu_{t})+\frac{1}{2}\Delta\nu_{t}\bigr{]}\Bigl{(}\bigl{(}f\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}-\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\frac{\mu_{t}}{\nu_{t}}\Bigr{)}\,\mathrm{d}x$ $=-\int\Bigl{\langle}b_{t}\nu_{t}-\frac{1}{2}\nabla\nu_{t},\nabla\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\frac{\mu_{t}}{\nu_{t}}\Bigr{\rangle}\,\mathrm{d}x$
@@
Lemma 20 (Functional inequalities under Gaussian convolutions, [chafai2004entropies, Corollary 13]).
@@
Theorem 21 ([chen2022improved, Theorem 3 and 4]).
@@
$\mathcal{R}_{q}(\mu P_{h}Q_{h}\mathbin{\|}\nu)\leq\begin{cases}\mathcal{R}_{q}(\mu\mathbin{\|}\nu)-\frac{2\log(1+h/C_{\mathsf{PI}})}{q}&\text{if }\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\geq 1\,,\\ \frac{\mathcal{R}_{q}(\mu\mathbin{\|}\nu)}{(1+h/C_{\mathsf{PI}})^{2}}&\text{if}\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\leq 1\,.\end{cases}$
@@
$\partial_{t}\mathcal{R}_{q}(\mu P_{t}\mathbin{\|}\nu P_{t})$ $=-\frac{q}{2}\,\frac{\mathbb{E}_{\mu P_{t}}\Bigl{[}\Bigl{\langle}\nabla\Bigl{(}\frac{\mu P_{t}}{\nu P_{t}}\Bigr{)}^{q-1},\nabla\log\frac{\mu P_{t}}{\nu P_{t}}\Bigr{\rangle}\Bigr{]}}{(q-1)\,\mathbb{E}_{\nu P_{t}}\Bigl{[}\Bigl{(}\frac{\mu P_{t}}{\nu P_{t}}\Bigr{)}^{q}\Bigr{]}}\underset{(i)}{=}-\frac{q}{2}\,\frac{\mathbb{E}_{\mu P_{t}}\Bigl{[}\Bigl{(}\frac{\mu P_{t}}{\nu P_{t}}\Bigr{)}^{q-2}\langle\nabla\frac{\mu P_{t}}{\nu P_{t}},\nabla\log\frac{\mu P_{t}}{\nu P_{t}}\rangle\Bigr{]}}{\mathbb{E}_{\nu P_{t}}\Bigl{[}\Bigl{(}\frac{\mu P_{t}}{\nu P_{t}}\Bigr{)}^{q}\Bigr{]}}$
=$\underset{(ii)}{=}-\frac{q}{2}\,\frac{\mathbb{E}_{\nu P_{t}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left(\frac{\mu P_{t}}{\nu P_{t}}}\right)^{q}\|\nabla\log\frac{\mu P_{t}}{\nu P_{t}}\|^{2}}\right]}{\mathbb{E}_{\nu P_{t}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left(\frac{\mu P_{t}}{\nu P_{t}}}\right)^{q}}\right]}=-\frac{1}{2}\,\mathsf{RF}_{q}(\mu P_{t}\mathbin{\|}\nu P_{t})\,,$
@@
$\partial_{t}\mathcal{R}_{q}(\mu P_{h}Q_{t}\mathbin{\|}\tilde{\nu}Q_{t})=-\frac{1}{2}\,\mathsf{RF}_{q}(\mu P_{h}Q_{t}\mathbin{\|}\tilde{\nu}Q_{t})\leq-\frac{1}{q}\,\frac{\mathcal{R}_{q}(\mu P_{h}Q_{t}\mathbin{\|}\tilde{\nu}Q_{t})}{C_{\mathsf{LSI}}(\tilde{\nu}Q_{t})}\underset{(i)}{\leq}-\frac{1}{q}\,\frac{\mathcal{R}_{q}(\mu P_{h}Q_{t}\mathbin{\|}\tilde{\nu}Q_{t})}{C_{\mathsf{LSI}}+{h-t}}\,,$
@@
The result in the $\mathcal{R}_{q}$ under (7) can be shown in a similar manner. Only difference is that in forward and backward computations, one should use the functional inequality in Lemma 16 and the following standard inequalities:
$1-\exp\bigl{(}-\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\bigr{)}\geq\begin{cases}\frac{1}{2}&\text{if }\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\geq 1\,,\\ \frac{1}{2}\mathcal{R}_{q}(\mu\mathbin{\|}\nu)&\text{if }\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\leq 1\,.\end{cases}$
@@
$\mu_{h}(x)=\frac{1}{(2\pi h)^{d/2}}\int\exp\bigl{(}-\frac{\|x-y\|^{2}}{2h}\bigr{)}\,\mathrm{d}\mu_{0}(y)\,.$

### v2

@@ 1.3 Discussion
Previous uniform samplers like $\mathsf{Ball\ walk}$ are made lazy (i.e., with probability $1/2$, it does nothing), to ensure convergence to the target stationary distribution. However, our algorithm does not need this, as our sampler is shown to directly contract towards the target.
@@
We remark that Theorem 5 places the previously known mixing guarantees for $\mathsf{Ball\ walk},\mathsf{Speedy\ walk}$ in a unified framework. Existing tight guarantees for $\mathsf{Speedy\ walk}$ are in TV distance and based on the log-Sobolev constant, assuming an oracle for implementing each step [LV17]. The known convergence guarantees of $\mathsf{Ball\ walk}$ (see §B for details), namely the mixing time of $\widetilde{\mathcal{O}}(Md^{2}D^{2}\log\frac{1}{\varepsilon})$ for TV distance, are for the composite algorithm [$\mathsf{Speedy\ walk}+$rejection sampling]. Here $\mathsf{Speedy\ walk}$ records only the accepted steps of $\mathsf{Ball\ walk}$, so its stationary distribution differs slightly from the uniform distribution (and can be corrected with a post-processing step). On the other hand, $\mathsf{In\text{-}and\text{-}Out}$ actually converges to $\pi^{\mathcal{K}}$ without any adjustments and achieves stronger Rényi divergence bounds in the same asymptotic complexity. Our analysis shows that the mixing guarantee is determined by isoperimetric constants of the target (Poincaré or log-Sobolev).
@@
The $\mathsf{Ball\ walk}$’s largest possible step size is of order $1/\sqrt{d}$ (see §B) to keep the rejection probability bounded by a constant. This bound could also be viewed as an “effective” step size of $\mathsf{In\text{-}and\text{-}Out}$, since the $\ell_{2}$-norm of the Gaussian $\mathcal{N}(0,hI)$ is concentrated around $\sqrt{hd}$ and we will set the variance $h$ of $\mathsf{In\text{-}and\text{-}Out}$ to $\widetilde{\mathcal{O}}(1/d^{2})$, so we have $\sqrt{hd}\approx 1/\sqrt{d}$.
@@
How can we compare $\mathsf{In\text{-}and\text{-}Out}$ with $\mathsf{Speedy\ walk}?$ Iterating speedy steps leads to a biased distribution. As clarified in Remark 3, one step of (a Gaussian version of) $\mathsf{Speedy\ walk}$ can be understood as a step of backward heat flow. Therefore, if one can control the isoperimetric constants of the biased distribution along the trajectory of the backward flow, then contraction of $\mathsf{Speedy\ walk}$ toward the biased distribution will follow from the simultaneous backward analysis.
@@ 2 Preliminaries
$\lVert\mu-\nu\rVert_{\mathsf{TV}}:=\sup_{B\in\mathcal{F}}\lvert\mu(B)-\nu(B)\rvert\,,$
@@
$\mathcal{W}_{2}^{2}(\mu,\nu):=\inf_{\gamma\in\Gamma(\mu,\nu)}\mathbb{E}_{(X,Y)\sim\gamma}[\lVert X-Y\rVert^{2}]\,,$
@@
We recall two important functional inequalities of a distribution.
@@
We say that a probability measure $\nu$ on $\mathbb{R}^{d}$ satisfies a Poincaré inequality (PI) with parameter $C_{\mathsf{PI}}(\nu)$ if for all smooth functions $f:\mathbb{R}^{d}\to\mathbb{R}$,
$\operatorname{var}_{\nu}f\leq C_{\mathsf{PI}}(\nu)\,\mathbb{E}_{\nu}[\lVert\nabla f\rVert^{2}]\,,$
where $\operatorname{var}_{\nu}f\coloneqq\mathbb{E}_{\nu}[\lvert f-\mathbb{E}_{\nu}f\rvert^{2}]$.
@@
We say that a probability measure $\nu$ on $\mathbb{R}^{d}$ satisfies a log-Sobolev inequality (LSI) with parameter $C_{\mathsf{LSI}}(\nu)$ if for all smooth functions $f:\mathbb{R}^{d}\to\mathbb{R}$,
$\mathsf{Ent}_{\nu}(f^{2})\leq 2C_{\mathsf{LSI}}(\nu)\,\mathbb{E}_{\nu}[\lVert\nabla f\rVert^{2}]\,,$
@@
where $\mathsf{FI}(\mu\mathbin{\|}\nu):=\mathbb{E}_{\mu}[\lVert\nabla\log\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\rVert^{2}]$ is the Fisher information of $\mu$ with respect to $\nu$.
We state two important lemmas which are needed for our proofs. The first is the data-processing inequality for Rényi divergence and $f$-divergence, given below.
@@
Functional inequalities allow us to show exponential contraction of various divergences, through the following helpful inequality.
@@ 3 Analysis
One successful iteration of $\mathsf{In\text{-}and\text{-}Out}$ is exactly the same as the composition of running the forward heat flow and then backward heat flow, both for time $h$.
@@
Regarding the backward part, it is known from [CCSW22, Lemma 14] that the construction of the time-reversal SDE ensures that $(Z_{h}^{\leftarrow},Z_{0}^{\leftarrow})$ and $(Z_{0},Z_{h})$ have the same joint distribution, when $Z_{0}\sim\pi^{X}$ (and so $Z_{h}\sim\pi^{Y}$). Hence, $\mathsf{law}(Z_{h}^{\leftarrow}|Z_{0}^{\leftarrow}=y)=\mathsf{law}(Z_{0}|Z_{h}=y)=\pi^{X|Y=y}$, where the last equality follows from $(Z_{0},Z_{h})\sim\pi$. Since we initialize ($\mathsf{BH}$) with $Z_{0}^{\leftarrow}=y\sim\mu_{k+1}^{Y}$, we have
@@
For (1), we collect in §3.1 some important implications of functional inequalities, e.g. the Poincaré and log-Sobolev inequalities, for the uniform distribution. Then in §3.2, we exploit the flow perspective of the algorithm to obtain the mixing guarantees. To this end, we revisit the proofs for the contraction results of forward and backward heat flows, and then extend them to our constrained setting:
Theorem 13 (Adapted for uniform distributions, [CCSW22]).
@@
The result reduces the problem of obtaining a mixing guarantee to that of demonstrating a functional inequality on the target distribution. For this, it is not strictly necessary that $\mathcal{K}$ be convex.
Convexity of $\mathcal{K}$ is crucial this time unlike (1). We show in §3.3 that the failure probability remains under control by taking a suitable variance $h$ and threshold $N$, and that the expected number of trials per iteration is of order $\log N$, not $N$.
@@
Let $\mathcal{K}$ be any convex body in $\mathbb{R}^{d}$ presented by a well-defined membership oracle, $\pi^{X}$ the uniform distribution over $\mathcal{K}$, and $\mu$ an $M$-warm initial distribution with respect to $\pi^{X}$. For any given $m\in\mathbb{N}$ and $\eta\in(0,1)$, set $Z=\frac{9mM}{\eta}(\geq 9)$, $h=\frac{\log\log Z}{2d^{2}\log Z}$ and $N=Z\log^{4}Z=\widetilde{\mathcal{O}}(\nicefrac{{mM}}{{\eta}})$. Then, the failure probability of one iteration of $\mathsf{In\text{-}and\text{-}Out}$ is at most $\eta/m$, and the expected membership queries per iteration is $\mathcal{O}\bigl{(}M\log^{4}\nicefrac{{mM}}{{\eta}}\bigr{)}$.
@@ 3.1 Functional inequalities
The contraction of an outer loop of our algorithm is controlled by isoperimetry of the uniform distribution $\pi^{X}$, which is described precisely by a functional inequality. The most natural ones to consider in this setting are the Poincaré inequality (7) and log-Sobolev inequality (8). In §C, we provide a more detailed discussion of how these are related to other important notions of isoperimetry, such as the Cheeger and log-Cheeger inequalities.
@@
Lemma 16 ([VW19, Lemma 9]).
@@
where $\mathsf{RF}_{q}(\mu\mathbin{\|}\nu):=q\,\mathbb{E}_{\nu}\bigl{[}\bigl{(}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\bigr{)}^{q}\lVert\nabla\log\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\rVert^{2}\bigr{]}\,/\,\mathbb{E}_{\nu}\bigl{[}\bigl{(}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\bigr{)}^{q}\bigr{]}$ is the Rényi Fisher information of order $q$ of $\mu$ with respect to $\nu$.
The log-Sobolev inequality paired with the KL-divergence (8) can be understood as a special case of the following inequality [^2] Such inequalities are often called Polyak-Łojasiewicz inequalities, which say for $f:\mathbb{R}^{d}\to\mathbb{R}$, and all $y\in\mathbb{R}^{d}$ that $f(y)\leq c\,\lVert\nabla f(y)\rVert^{2}$ for some constant $c$, if $\min f(x)=0$. paired with the $q$-Rényi divergence for $q\geq 1$.
Lemma 17 ([VW19, Lemma 5]).
@@
We have collected below the functional inequalities used to establish the mixing of our algorithm (see §C for a detailed presentation).
@@
Let $\mathcal{K}\subset\mathbb{R}^{d}$ be a convex body with diameter $D$, and $\pi$ be the uniform distribution over $\mathcal{K}$. Then, $C_{\mathsf{PI}}(\pi)\lesssim\lVert\operatorname{cov}\pi\rVert_{\mathsf{op}}\,\log d$ and $C_{\mathsf{LSI}}(\pi)\lesssim D^{2}$. If $\pi$ is isotropic, then $C_{\mathsf{PI}}(\pi)\lesssim\log d$ and $C_{\mathsf{LSI}}(\pi)\lesssim D$.
@@ 3.2 Contraction and mixing
We start by analyzing how many outer iterations of $\mathsf{In\text{-}and\text{-}Out}$ are required to be $\varepsilon$-close to $\pi^{X}$, the uniform distribution over $\mathcal{K}$. The contraction of Algorithm 1 comes from analyzing Lines 2 and 3 through the perspective of heat flows (see §1.1). To exploit this view, we first revisit the previous contraction analysis in [CCSW22], which is carried out for distributions with smooth densities. Although the uniform distribution is not even continuous, we prove a technical lemma (Lemma 22) that enables us to extend previously known results to the uniform distribution. Lastly, combining the previous results with our technical lemma, we obtain clean contraction results of Algorithm 1 toward the uniform distribution $\pi^{X}$ in Theorem 23.
@@
In this part, we review the contraction results for heat flow and its time-reversal [CCSW22], which are intimately connected with our algorithm. We also provide key technical ingredients needed for its proof, such as the computations for measures evolving under simultaneous forward/backward heat flows. We refer interested readers to §D for additional details. Only in Part I, we assume that $\nu$ denotes a probability measure with smooth density.
@@
Although there are many ways to define a “reversal” of $P_{h}$, we will use the notion of adjoint introduced by [KP21], which is the most immediately useful.
@@
Instead of considering the change in metrics along the evolution of $\mu P_{t}$ with respect to “fixed” $\nu$, it will be useful to consider the simultaneous evolution of $\mu P_{t},\nu P_{t}$ (and similarly $\mu Q_{t},(\nu P_{h})Q_{t}$). This type of computation was carried out for specific metrics in earlier work [VW19, CCSW22]. The following is a more generalized form of one appearing in [YFL+23, Lemma 2]. In the lemma below, we consider an arbitrary diffusion equation with corresponding Fokker-Planck equation:
@@
$\int\bigl{[}-\nabla\cdot(b_{t}\nu_{t})+$ $\frac{1}{2}\Delta\nu_{t}\bigr{]}\Bigl{(}\bigl{(}f\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}-\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\frac{\mu_{t}}{\nu_{t}}\Bigr{)}\,\mathrm{d}x$
$=-\int\Bigl{\langle}b_{t}\nu_{t}-\frac{1}{2}\nabla\nu_{t},\nabla\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\frac{\mu_{t}}{\nu_{t}}\Bigr{\rangle}\,\mathrm{d}x$
@@
Lemma 20 (Functional inequalities under Gaussian convolutions, [Cha04, Corollary 13]).
@@
Theorem 21 ([CCSW22, Theorem 3 and 4]).
@@
$\mathcal{R}_{q}(\mu P_{h}Q_{h}\mathbin{\|}\nu)\leq\begin{cases}\mathcal{R}_{q}(\mu\mathbin{\|}\nu)-\frac{2\log(1+h/C_{\mathsf{PI}})}{q}&\text{if }\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\geq 1\,,\\ \frac{\mathcal{R}_{q}(\mu\mathbin{\|}\nu)}{(1+h/C_{\mathsf{PI}})^{2/q}}&\text{if }\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\leq 1\,.\end{cases}$
@@
$\partial_{t}\mathcal{R}_{q}(\mu P_{t}\mathbin{\|}\nu P_{t})$ $=-\frac{q}{2}\,\frac{\mathbb{E}_{\mu P_{t}}\Bigl{[}\Bigl{\langle}\nabla\Bigl{(}\frac{\mu P_{t}}{\nu P_{t}}\Bigr{)}^{q-1},\nabla\log\frac{\mu P_{t}}{\nu P_{t}}\Bigr{\rangle}\Bigr{]}}{(q-1)\,\mathbb{E}_{\nu P_{t}}\Bigl{[}\Bigl{(}\frac{\mu P_{t}}{\nu P_{t}}\Bigr{)}^{q}\Bigr{]}}$
=$\underset{(i)}{=}-\frac{q}{2}\,\frac{\mathbb{E}_{\mu P_{t}}\Bigl{[}\Bigl{(}\frac{\mu P_{t}}{\nu P_{t}}\Bigr{)}^{q-2}\langle\nabla\frac{\mu P_{t}}{\nu P_{t}},\nabla\log\frac{\mu P_{t}}{\nu P_{t}}\rangle\Bigr{]}}{\mathbb{E}_{\nu P_{t}}\Bigl{[}\Bigl{(}\frac{\mu P_{t}}{\nu P_{t}}\Bigr{)}^{q}\Bigr{]}}$
=$\underset{(ii)}{=}-\frac{q}{2}\,\frac{\mathbb{E}_{\nu P_{t}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left(\frac{\mu P_{t}}{\nu P_{t}}}\right)^{q}\lVert\nabla\log\frac{\mu P_{t}}{\nu P_{t}}\rVert^{2}}\right]}{\mathbb{E}_{\nu P_{t}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left(\frac{\mu P_{t}}{\nu P_{t}}}\right)^{q}}\right]}=-\frac{1}{2}\,\mathsf{RF}_{q}(\mu P_{t}\mathbin{\|}\nu P_{t})\,,$
@@
$\partial_{t}\mathcal{R}_{q}(\mu P_{h}Q_{t}\mathbin{\|}\tilde{\nu}Q_{t})$ $=-\frac{1}{2}\,\mathsf{RF}_{q}(\mu P_{h}Q_{t}\mathbin{\|}\tilde{\nu}Q_{t})$
$\leq-\frac{1}{q}\,\frac{\mathcal{R}_{q}(\mu P_{h}Q_{t}\mathbin{\|}\tilde{\nu}Q_{t})}{C_{\mathsf{LSI}}(\tilde{\nu}Q_{t})}\underset{(i)}{\leq}-\frac{1}{q}\,\frac{\mathcal{R}_{q}(\mu P_{h}Q_{t}\mathbin{\|}\tilde{\nu}Q_{t})}{C_{\mathsf{LSI}}+{h-t}}\,,$
@@
The result in the $\mathcal{R}_{q}$ under (7) can be shown in a similar manner. The only difference is that in forward and backward computations, one should use the functional inequality in Lemma 16 and the following standard inequalities:
$1-\exp\bigl{(}-\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\bigr{)}\geq\begin{cases}\frac{1}{2}&\text{if }\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\geq 1\,,\\ \frac{1}{2}\mathcal{R}_{q}(\mu\mathbin{\|}\nu)&\text{if }\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\leq 1\,.\end{cases}\qed$
@@
$\mu_{h}(x)=\frac{1}{(2\pi h)^{d/2}}\int\exp\bigl{(}-\frac{\lVert x-y\rVert^{2}}{2h}\bigr{)}\,\mathrm{d}\mu_{0}(y)\,.$